# Pullback of arithmetic theta series and its modularity for unitary Shimura curves

### v1

@@ 5. Pullback of algebraic divisors
Assume $m\in\mathbb{Z}$. Then
(5.8) $\mathcal{Z}^{\diamond,*}(m)|_{\mathcal{S}}\cong\displaystyle\bigotimes_{\begin{subarray}{c}m_{1}+m_{2}=m\\ m_{1}\neq 0\end{subarray}}\mathcal{Z}^{*}(m_{1})^{\otimes r_{\Lambda}(m_{2})}\bigotimes(\Omega^{-1})^{\otimes r_{\Lambda}(m)},$
@@
The proof is similar to that of [HMP20, Proposition 6.6.3]. If $m<0$, then both sides of (5.8) are trivial line bundles. If $m=0$, then (5.8) is simply (5.5) by §3.2. For any scheme $S\rightarrow\mathcal{S}$, let $\mathbf{A}_{S}$ be the pullback of the universal abelian scheme over $\mathcal{S}$, and $\mathbf{A}_{S}^{\diamond}$ be the pullback of the universal abelian scheme over $\mathcal{S}^{\diamond}$ via the morphism $S\rightarrow\mathcal{S}\xrightarrow[]{\varphi_{\Lambda}}\mathcal{S}^{\diamond}$. Define the following $\mathcal{O}_{\boldsymbol{k}}$-modules of special homomorphisms
$V(\mathbf{A}_{S})=\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}(\mathbf{A}_{0,S},\mathbf{A}_{S}),\quad V(\mathbf{A}_{S}^{\diamond})=\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}(\mathbf{A}_{0,S},\mathbf{A}_{S}^{\diamond}).$
By the definition of $\varphi_{\Lambda}$, we know $\mathbf{A}_{S}^{\diamond}=\mathbf{A}_{S}^{\diamond}\oplus(\mathbf{A}_{0}\otimes_{\mathcal{O}_{\boldsymbol{k}}}\Lambda)_{S}$. By [AK18, Proposition 2], we know that $\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}(\mathbf{A}_{0},\mathbf{A}_{0}\otimes_{\mathcal{O}_{\boldsymbol{k}}}\Lambda)\cong\Lambda$. Hence we have
(5.9) $V(\mathbf{A}_{S}^{\diamond})\cong V(\mathbf{A}_{S})\oplus\Lambda.$
Equation (5.9) induces the following isomorphism of $\mathcal{S}$-stacks.
$\mathcal{Z}^{\diamond}(m)|_{\mathcal{S}}\cong\bigsqcup_{\begin{subarray}{c}m_{1}+m_{2}=m\\ m_{1}\neq 0\\ \lambda\in R_{\Lambda}(m_{2})\end{subarray}}\mathcal{Z}(m_{1})\sqcup\bigsqcup_{\lambda\in R_{\Lambda}(m)}\mathcal{S}.$
By taking Zariski closure we have the the following isomorphism of $\mathcal{S}^{*}$-stacks.
@@
Let $\mathcal{Z}^{\diamond}$ be a connected component of $\mathcal{Z}^{\diamond}_{\mathrm{im}}$ and $U^{\diamond}$ a connected étale neighborhood of $\mathcal{S}^{\diamond,*}$ such that $\mathcal{Z}^{\diamond}|_{U^{\diamond}}\rightarrow U^{\diamond}$ is a closed immersion. Let $U$ be a connected étale neighborhood of $\mathcal{S}^{*}$ whose image under $\varphi_{\Lambda}$ is contained in $\mathcal{Z}^{\diamond}|_{U^{\diamond}}$. Then we have the following equality of line bundles.
$\mathcal{O}(\mathcal{Z}^{\diamond})|_{U}\cong\Omega^{-1}|_{U}.$
@@
The proof is identical to that of [BHY15, Theorem 7.10] or [HMP20, Lemma 6.6.4], given Theorem 4.4. We leave it to the reader. ∎
@@
(5.12) $\mathcal{B}(m,v)\coloneqq\frac{1}{4\pi v}\sum_{\Phi\in\mathrm{Cusp}([[L]])}\#\left\{x\in L_{\Phi}:\langle x,x\rangle=m\right\}\cdot\mathcal{S}^{*}(\Phi)\in\mathrm{CH}^{1}_{\mathbb{C}}(\mathcal{S}^{*}),$
@@
$=$ $\frac{1}{4\pi v}\sum_{\Phi^{\diamond}\in\mathrm{Cusp}([[L^{\diamond}]])}\#\left\{x\in L_{\Phi^{\diamond}}:\langle x,x\rangle=m\right\}\cdot\varphi_{\Lambda}^{*}(\mathcal{S}^{\diamond,*}(\Phi^{\diamond}))$
$=$ $\frac{1}{4\pi v}\sum_{\Phi\in\mathrm{Cusp}([[L]])}\#\left\{x\in L_{\Phi}\obot\Lambda:\langle x,x\rangle=m\right\}\cdot\mathcal{S}^{*}(\Phi)$
$=$ $\frac{1}{4\pi v}\sum_{\Phi\in\mathrm{Cusp}([[L]])}\sum_{m_{1}+m_{2}=m}\#\left\{x\in\Lambda:\langle x,x\rangle=m_{2}\right\}\cdot\#\left\{x\in L_{\Phi}:\langle x,x\rangle=m_{1}\right\}\cdot\mathcal{S}^{*}(\Phi)$
$=$ $\sum_{m_{1}+m_{2}=m}\#\left\{x\in\Lambda:\langle x,x\rangle=m_{2}\right\}\cdot\frac{1}{4\pi v}\sum_{\Phi\in\mathrm{Cusp}([[L]])}\#\left\{x\in L_{\Phi}:\langle x,x\rangle=m_{1}\right\}\cdot\mathcal{S}^{*}(\Phi)$
@@
6. Pullback of arithmetic divisors
@@ 6.1. Arithmetic Chow groups
Let $M^{*}$ be a compact complex manifold of dimension $n-1,\partial M^{*}\subset$ $M^{*}$ be a smooth codimension one submanifold, $M=M^{*}\backslash\partial M^{*}$, and $z_{0}\in\partial M^{*}$. Suppose that on some open neighborhood $V\subset M^{*}$ of $z_{0}$, there are coordinates $q,u_{1},\ldots,u_{n-2}$ such that $\partial M^{*}$ is the vanishing locus of $q$. After possibily shrinking $V$, we may always assume that $\log\left|q^{-1}\right|>1$ on $V$. Then we call the open set $V$ and its coordinates adapted to $\partial M^{*}$.
@@
Let $\mathcal{Z}=\sum m_{i}\mathcal{Z}_{i}$ be a finite $\mathbb{R}$-linear combination of pairwise distinct irreducible closed substacks of codimension one where $m_{i}\in\mathbb{R}$, which is a divisor on $\mathcal{S}^{*}$ with real coefficients.
@@
A Green function for $\mathcal{Z}$ consists of a smooth real-valued function $\operatorname{Gr}(\mathcal{Z},\cdot)$ on $\mathcal{S}(\mathbb{C})\backslash\mathcal{Z}(\mathbb{C})$ satisfying the following properties:
@@
We define the first arithmetic Chow group $\widehat{\mathrm{CH}}_{\mathbb{C}}^{1}\left(\mathcal{S}^{*}\right)$ to be the quotient
$\mathrm{Span}_{\mathbb{C}}\{\text{arithmetic divisors}\}/\mathrm{Span}_{\mathbb{C}}\{\text{principal arithmetic divisors}\}.$
Accordingly, we define $\widehat{\mathrm{Pic}}_{\mathbb{C}}(\mathcal{S}^{*})$ to be the isomorphism classes of metrized line bundles $\widehat{\mathcal{L}}=(\mathcal{L},||\cdot||)$ such that there exists a section $s$ of $\mathcal{L}$ with $-\log||s||^{2}=g$ on $\mathcal{S}(\mathbb{C})$. Here $g$ is a Green function (in the sense of Definition 6.3) for the divisor corresponding to $\mathcal{L}$. Then we have a natural isomorphism
(6.2) $i:\widehat{\mathrm{Pic}}_{\mathbb{C}}(\mathcal{S}^{*})\xrightarrow{\sim}\widehat{\mathrm{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*})$
such that under $i$, an element $\widehat{\mathcal{L}}=(\mathcal{L},||\cdot||)$ goes to the class of $(\hbox{div}(s),-\log||s||^{2})$ for some non-zero meromorphic section $s$ of $\mathcal{L}$. Conversely, the preimage of the class of $(\mathcal{Z},g)$ under $i$ is represented by
@@
where $-\log||{\bf 1}||^{2}=g$ with ${\bf 1}$ the canonical section of $\mathcal{O}(\mathcal{Z})$ and $g$ a Green function for $\mathcal{Z}$. In particular, we can define the pullback of arithmetic divisors to be the pullback of the corresponding metrized line bundles.
6.2. Arithmetic divisors
@@
Then $\mathcal{D}(x)$ is nonempty if and only if $(x,x)>0$. Also define
$\Gamma_{j,x}=\{\gamma\in\Gamma_{j}\mid\gamma x=x\},$
and $Z(x)_{j}$ to be the image of the natural complex immersion $\Gamma_{j,x}\backslash\mathcal{D}(x)\rightarrow\Gamma_{j}\backslash\mathcal{D}$.
@@
$Z(m)=\bigsqcup_{\mathrm{Cl}(\boldsymbol{k})}\bigsqcup_{L_{j}\in[[L]]}\bigsqcup_{\begin{subarray}{c}x\in L_{j}\\ (x,x)=m\end{subarray}}Z(x)_{j}.$
@@
where $\gamma$ is the Euler’s constant. In particular, $\beta_{1}(r)$ has log singularity at $r=0$. For $x\in L_{j}$ and $v>0$, we define the Kudla’s Green function on $\Gamma_{j}\backslash D$
@@
Then $\mathrm{Gr}(m,v)$ is a Green function for $Z(m)$. Note that $Z(m)=0$ and $\mathrm{Gr}(m,v)$ is smooth on $\mathcal{D}$ for $m\leq 0$. The following theorem describes the behavior of $\mathrm{Gr}(m,v)$ near the boundary.
@@
$\mathcal{E}(z)=\operatorname{Gr}(m,v)(z)+\log\left|\psi_{m}(z)\right|^{2}+\frac{\#\left\{x\in L_{\Phi}:\langle x,x\rangle=m\right\}}{4\pi v}\log|q(z)|^{2}$
@@
(6.9) $\widehat{\mathcal{Z}}_{K}^{\mathrm{tot}}(m,v)\coloneqq(\mathcal{Z}^{\mathrm{tot}}(m,v),\mathrm{Gr}(m,v))\in\widehat{\operatorname{CH}}^{1}(\mathcal{S}^{*}).$
In the rest of this section, we often drop the subscript $K$ for simplicity.
Now we deal with the case $m=0$. Recall that on $\mathcal{S}(\mathbb{C})$ the two line bundles $\Omega$ and $\omega$ agree, and can be identified the line bundle (still denoted by $\Omega$) of modular forms on $\mathcal{S}(\mathbb{C})$ by [BHY15, Proposition 6.2]. Moreover, this line bundle is descended from the tautological line bundle on $\mathcal{D}$, again denoted by $\Omega$. Let $s$ be a section of $\Omega$ and for $z\in\mathcal{D}$, we let $s_{z}\in z\subset V_{\mathbb{R}}$ denote the value of $s$ at $z$. Then we can define a metric on $\Omega$ by setting
@@
Here $\gamma$ is the Euler constant and $v>0$ is an extra parameter (eventually imaginary part of modular variable $\tau\in\mathbb{H}$). This gives an element $\widehat{\Omega}^{-1}\in\widehat{\hbox{Pic}}_{\mathbb{C}}(\mathcal{S})$. By [BHY15, Proposition 6.3], this element extends naturally to an element, still denoted by $\widehat{\Omega}^{-1}$ in $\widehat{\mathrm{Pic}}_{\mathbb{C}}(\mathcal{S}^{*})\cong\widehat{\mathrm{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*})$.
Following [ES18, Page 34], We define $\widehat{\mathcal{Z}}^{\mathrm{tot}}(0,v)$ to be
(6.11) $\widehat{\mathcal{Z}}^{\mathrm{tot}}(0,v)=\widehat{\Omega}^{-1}+(\mathcal{B}(0,v),\mathrm{Gr}(0,v))\in\widehat{\mathrm{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*}),$
where $\mathcal{B}(0,v)$ is defined in (5.12). Although the metrized line bundle $\widehat{\Omega}^{-1}$ depends on the number $v>0$, we omit it from the notation for convenience.
@@
As in §5, we use the superscript $\diamond$ to denote objects on $\mathcal{S}^{\diamond}$ or $\mathcal{S}^{\diamond,*}$. For each $L_{j}\in[[L]]$, define $L_{j}^{\diamond}=L_{j}\obot\Lambda$. The aim of this subsection is to prove Theorem 1.2, which is equivalent to the following theorem.
@@
Assume $m\in\mathbb{Z}$ and $v>0$. Then we have the following equation in $\widehat{\mathrm{CH}}^{1}_{\mathbb{C}}(\mathcal{S}^{*})$.
(6.12) $\varphi_{\Lambda}^{*}(\widehat{\mathcal{Z}}^{\diamond,\mathrm{tot}}(m,v))=\sum_{m_{1}+m_{2}=m}r_{\Lambda}(m_{2})\cdot\widehat{\mathcal{Z}}^{\mathrm{tot}}(m_{1},v).$

### v2

@@ 5.3. Pullback of line bundles
Assume $m\in\mathbb{Z}$. Then in $\hbox{Pic}(\mathcal{S}^{*})$
(5.19) $\mathcal{Z}^{\diamond,*}(m)|_{\mathcal{S}^{*}}\cong\displaystyle(\bigotimes_{\begin{subarray}{c}m_{1}+m_{2}=m\\ m_{1}\neq 0\end{subarray}}\mathcal{Z}^{*}(m_{1})^{\otimes r_{\Lambda}(m_{2})})\bigotimes(\Omega^{-1})^{\otimes r_{\Lambda}(m)},$
@@
The proof is similar to that of [HMP20, Proposition 6.6.3]. If $m<0$, then both sides of (5.19) are trivial line bundles. If $m=0$, then (5.19) is simply (6.14) by the definition of $\mathcal{Z}^{*}(0)$ in §3.2. For any scheme $S\rightarrow\mathcal{S}$, let $\mathbf{G}_{S}$ be the pullback of the universal semi-abelian scheme over $\mathcal{S}$, and $\mathbf{G}_{S}^{\diamond}$ be the pullback of the universal semi-abelian scheme over $\mathcal{S}^{\diamond}$ via the morphism $S\rightarrow\mathcal{S}\xrightarrow[]{\varphi_{\Lambda}}\mathcal{S}^{\diamond}$. Define the following $\mathcal{O}_{\boldsymbol{k}}$-modules of special homomorphisms
$V(\mathbf{G}_{S})=\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}(\mathbf{A}_{0,S},\mathbf{G}_{S}),\quad V(\mathbf{G}_{S}^{\diamond})=\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}(\mathbf{A}_{0,S},\mathbf{G}_{S}^{\diamond}).$
By the definition of $\varphi_{\Lambda}$, we know $\mathbf{G}_{S}^{\diamond}=\mathbf{G}_{S}\oplus(\mathbf{A}_{0}\otimes_{\mathcal{O}_{\boldsymbol{k}}}\Lambda)_{S}$. By [AK18, Proposition 2], we know that $\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}(\mathbf{A}_{0},\mathbf{A}_{0}\otimes_{\mathcal{O}_{\boldsymbol{k}}}\Lambda)\cong\Lambda$. Hence we have
(5.20) $V(\mathbf{G}_{S}^{\diamond})\cong V(\mathbf{G}_{S})\oplus\Lambda.$
Equation (5.20) induces the following isomorphism of $\mathcal{S}^{*}$-stacks.
@@
Let $\mathcal{Z}^{\diamond}$ be a connected component of $\mathcal{Z}^{\diamond}_{\mathrm{im}}$ and $U^{\diamond}$ a connected étale neighborhood of $\mathcal{S}^{\diamond,*}$ such that the morphism $\mathcal{Z}^{\diamond}|_{U^{\diamond}}\rightarrow U^{\diamond}$ is a closed immersion when restricted on each connected component of $\mathcal{Z}^{\diamond}|_{U^{\diamond}}$. Let $U$ be a connected étale neighborhood of $\mathcal{S}^{*}$ whose image under $\varphi_{\Lambda}$ is contained in $\mathcal{Z}^{\diamond}|_{U^{\diamond}}$. Then we have canonical isomorphism of line bundles
$\mathcal{Z}^{\diamond}_{\mathrm{prop}}=\displaystyle\bigotimes_{\begin{subarray}{c}m_{1}+m_{2}=m\\ m_{1}\neq 0\end{subarray}}\mathcal{Z}^{*}(m_{1})|_{U}^{\otimes r_{\Lambda}(m_{2})},$
and
(5.23) $\mathcal{O}(\mathcal{Z}^{\diamond})|_{U}\cong\Omega^{-1}|_{U}.$
@@
The proof is identical to that of [BHY15, Theorem 7.10] or [HMP20, Lemma 6.6.4], given Proposition 4.1 and Proposition 4.8. We leave it to the reader. ∎
@@
(5.24) $\mathcal{B}(m,v)\coloneqq\frac{1}{4\pi v}\sum_{\Phi\in\mathrm{Cusp}([[L]])}\#\left\{x\in L_{\Phi}:(x,x)=m\right\}\cdot\mathcal{S}^{*}(\Phi)\in\mathrm{CH}^{1}_{\mathbb{C}}(\mathcal{S}^{*}),$
@@
$=$ $\frac{1}{4\pi v}\sum_{\Phi^{\diamond}\in\mathrm{Cusp}([[L^{\diamond}]])}\#\left\{x\in L_{\Phi^{\diamond}}:(x,x)=m\right\}\cdot\varphi_{\Lambda}^{*}(\mathcal{S}^{\diamond,*}(\Phi^{\diamond}))$
$=$ $\frac{1}{4\pi v}\sum_{\Phi\in\mathrm{Cusp}([[L]])}\#\left\{x\in L_{\Phi}\obot\Lambda:(x,x)=m\right\}\cdot\mathcal{S}^{*}(\Phi)$
$=$ $\frac{1}{4\pi v}\sum_{\Phi\in\mathrm{Cusp}([[L]])}\sum_{m_{1}+m_{2}=m}\#\left\{x\in\Lambda:(x,x)=m_{2}\right\}\cdot\#\left\{x\in L_{\Phi}:(x,x)=m_{1}\right\}\cdot\mathcal{S}^{*}(\Phi)$
$=$ $\sum_{m_{1}+m_{2}=m}\#\left\{x\in\Lambda:(x,x)=m_{2}\right\}\cdot\frac{1}{4\pi v}\sum_{\Phi\in\mathrm{Cusp}([[L]])}\#\left\{x\in L_{\Phi}:(x,x)=m_{1}\right\}\cdot\mathcal{S}^{*}(\Phi)$
@@
6. Pullback of arithmetic special divisors
@@ 6.1. Arithmetic Chow groups
Let $M^{*}$ be a compact complex manifold of dimension $n-1,\partial M^{*}\subset$ $M^{*}$ be a smooth codimension one submanifold, $\MM=M^{*}\backslash\partial M^{*}$, and $z_{0}\in\partial M^{*}$. Suppose that on some open neighborhood $V\subset M^{*}$ of $z_{0}$, there are coordinates $q,u_{1},\ldots,u_{n-2}$ such that $\partial M^{*}$ is the vanishing locus of $q$. After possibily shrinking $V$, we may always assume that $\log\left|q^{-1}\right|>1$ on $V$. Then we call the open set $V$ and its coordinates adapted to $\partial M^{*}$.
@@
Let $\mathcal{Z}=\sum m_{i}\mathcal{Z}_{i}$ be a finite $\mathbb{C}$-linear combination of pairwise distinct irreducible closed substacks of codimension one where $m_{i}\in\mathbb{C}$, which is a divisor on $\mathcal{S}^{*}$ with complex coefficients.
@@
A Green function for $\mathcal{Z}$ consists of a smooth function $\operatorname{Gr}(\mathcal{Z},\cdot)$ on $\mathcal{S}(\mathbb{C})\backslash\mathcal{Z}(\mathbb{C})$ satisfying the following properties:
@@
We recall the first arithmetic Chow group
$\widehat{\mathrm{CH}}^{1}\left(\mathcal{S}^{*}\right)=\mathrm{Span}\{\text{arithmetic divisors}\}/\mathrm{Span}\{\text{principal arithmetic divisors}\}.$
[…]
(6.2) $i:\widehat{\mathrm{Pic}}(\mathcal{S}^{*})\rightarrow\widehat{\mathrm{CH}}^{1}(\mathcal{S}^{*}).$
The preimage of the class of $(\mathcal{Z},g)$ under $i$ is represented by
@@
[…]
(6.3) $\widehat{\mathrm{Pic}}_{\mathbb{R}}(\mathcal{S}^{*})\cong\widehat{\mathrm{CH}}_{\mathbb{R}}^{1}(\mathcal{S}^{*}).$
On the Picard group side, we allow formally real power of a line bundle $\mathcal{L}^{r}$ with metric being a $r$-power of a metric on $\mathcal{L}$. On the arithmetic side, we allow the $r$-multiple of an honest divisor. We need this to accommodate the arithmetic divisors $\widehat{\mathcal{Z}}^{\mathrm{tot}}(m,v)$.
Note that according to [BGKK07, Proposition 7.5], we can define a pullback $\varphi_{\Lambda}^{*}:\widehat{\mathrm{Pic}}_{\mathbb{R}}(\mathcal{S}^{\diamond,*})\to\widehat{\mathrm{Pic}}_{\mathbb{R}}(\mathcal{S}^{*})$ since $\varphi_{\Lambda}^{-1}(\partial\mathcal{S}^{\diamond,*})\subset\varphi_{\Lambda}^{-1}(\partial\mathcal{S}^{*})$.
6.2. Arithmetic special divisors
@@
Then $\mathcal{D}(x)$ is nonempty if and only if $(x,x)>0$. Let $Z(x)_{j}$ to be the image of $\mathcal{D}(x)$ in $\Gamma_{j}\backslash\mathcal{D}$.
@@
$Z(m)=\bigsqcup_{\mathrm{Cl}(\boldsymbol{k})}\bigsqcup_{L_{j}\in[[L]]}\bigsqcup_{\begin{subarray}{c}x\in L_{j}\mod\Gamma_{j}\\ (x,x)=m\end{subarray}}Z(x)_{j}.$
@@
where $\gamma$ is the Euler’s constant. In particular, $\beta_{1}(r)$ has log singularity at $r=0$. For $0\neq x\in L_{j}$ and $v>0$, we define the Kudla’s Green function on $\D\Gamma_{j}\backslash$ D
@@
Then $\mathrm{Gr}(m,v)$ is a Green function for $Z(m)$. Note that $Z(m)=0$ and $\mathrm{Gr}(m,v)$ is smooth on $\mathcal{D}$ for $m\leq 0$. The following theorem states that $\mathrm{Gr}(m,v)$ is a Green function of $\mathcal{Z}^{\mathrm{tot}}(m).$
@@
$\mathcal{E}(z)=\operatorname{Gr}(m,v)(z)+\log\left|\psi_{m}(z)\right|^{2}+\frac{\#\left\{x\in L_{\Phi}:(x,x)=m\right\}}{4\pi v}\log|q(z)|^{2}$
@@
(6.10) $\widehat{\mathcal{Z}}^{\mathrm{tot}}(m,v)\coloneqq(\mathcal{Z}^{\mathrm{tot}}(m,v),\mathrm{Gr}(m,v))\in\widehat{\operatorname{CH}}_{\mathbb{R}}^{1}(\mathcal{S}^{*}).$
It is an arithmetic divisor with the Kudla Green function, denoted by $\widehat{\mathcal{Z}}_{K}^{\mathrm{tot}}(m,v)$ in the introduction. We will add subscript $K$ when needing to distinguish it from the arithmetic divisors $\widehat{\mathcal{Z}}_{B}^{\mathrm{tot}}(m,v)$ with the Bruinier Green functions
Now we deal with the case $m=0$. Recall that on $\mathcal{S}(\mathbb{C})$ the two line bundles $\Omega$ and $\omega$ agree, and can be identified the line bundle (still denoted by $\Omega$) of modular forms of weight $1$ on $\mathcal{S}(\mathbb{C})$ by [BHY15, Proposition 6.2]. Moreover, this line bundle is descended from the tautological line bundle on $\mathcal{D}$, again denoted by $\Omega$. Let $s$ be a section of $\Omega$ and for $z\in\mathcal{D}$, we let $s_{z}\in z\subset V_{\mathbb{R}}$ denote the value of $s$ at $z$. Then we can define a metric on $\Omega$ by setting
@@
Here $\gamma$ is the Euler constant and $v>0$ is an extra parameter (eventually imaginary part of modular variable $\tau\in\mathbb{H}$). Compared to the nomalization in [ES18, Page 34], we add a $v$-factor to count the arithmetic divisor $(0,\log v)$ there. This gives an element $\widehat{\Omega}^{-1}\in\widehat{\hbox{Pic}}(\mathcal{S})$ that depends on an extra factor $v$. By [BHY15, Proposition 6.3], this element extends naturally to an element, still denoted by $\widehat{\Omega}^{-1}$ in $\widehat{\mathrm{Pic}}(\mathcal{S}^{*})\hookrightarrow\widehat{\mathrm{CH}}_{\mathbb{R}}^{1}(\mathcal{S}^{*})$.
Following [ES18, Page 34], we define $\widehat{\mathcal{Z}}^{\mathrm{tot}}(0,v)$ as
(6.12) $\widehat{\mathcal{Z}}^{\mathrm{tot}}(0,v)=\widehat{\Omega}^{-1}+(\mathcal{B}(0,v),\mathrm{Gr}(0,v))\in\widehat{\mathrm{CH}}_{\mathbb{R}}^{1}(\mathcal{S}^{*}),$
where $\mathcal{B}(0,v)$ is defined in (5.24).
@@
As in §5, we use the superscript $\diamond$ to denote objects on $\mathcal{S}^{\diamond}$ or $\mathcal{S}^{\diamond,*}$. The aim of this subsection is to prove Theorem 1.3, which is equivalent to the following theorem.
@@
Assume $m\in\mathbb{Z}$ and $v>0$. Then the following equation
(6.13) $\varphi_{\Lambda}^{*}(\widehat{\mathcal{Z}}^{\diamond,\mathrm{tot}}(m,v))=\sum_{m_{1}+m_{2}=m}r_{\Lambda}(m_{2})\cdot\widehat{\mathcal{Z}}^{\mathrm{tot}}(m_{1},v)$
holds in $\widehat{\mathrm{CH}}^{1}_{\mathbb{R}}(\mathcal{S}^{*})$.
The right-hand side is an infinite sum and needs a little explanation as $m_{1}$ can be negative. It is equal to
$r_{\Lambda}(m)\widehat{\mathcal{Z}}(0,\mu)+(\sum_{\begin{subarray}{c}m_{1}+m_{2}=m\\ m_{1}>0\end{subarray}}r_{\Lambda}(m_{2})\mathcal{Z}^{\mathrm{tot}}(m_{1},v),\sum_{\begin{subarray}{c}m_{1}+m_{2}=m\\ m_{1}\neq 0\end{subarray}}r_{\Lambda}(m_{2})G(m_{1},v)).$
The sum over cycles is finite as $\mathcal{Z}^{\mathrm{tot}}(m_{1},v)=0$ for $m_{1}<0$, and the sum of Green functions is a convergent sum.